# Each Rank Could be an Expert: Single-Ranked Mixture of Experts LoRA for Multi-task Learning

### v1

@@ 4 Experiments
In Fig.6, we select a specific layer and illustrate the load distribution across all ranks, both with and without the auxiliary load balancing strategy. When the load balancing strategy is applied, the load is evenly distributed across ranks, indicating a relatively balanced workload. In contrast, without the load-balancing strategy, certain ranks experience significantly higher loads, leading to an imbalanced workload distribution. To quantitatively evaluate the load balance, we calculate the MaxVio (maximal violation) value (Wang et al., 2024), defined as the difference between the maximum load and the average load, with smaller values indicating a more balanced distribution. With the load balancing strategy, the MaxVio is significantly reduced to 1.23, compared to 5.71 without the strategy. This stark contrast further highlights the effectiveness of the adopted load-balancing strategy in achieving a more uniform load distribution. As shown in the right part of Fig.6, applying the load-balancing strategy results in a 0.34% performance improvement. In Appendix G, we further visualize the routing distribution under the ablation of the load-balancing strategy. The results clearly demonstrate that with the load-balancing strategy removed, the routing becomes unevenly distributed, heavily favoring a subset of experts. This imbalance highlights the critical role of the load-balancing mechanism in ensuring a fair and efficient allocation of tasks across all available experts.
@@
In our approach, the router selects the appropriate ranks to participate in the computation for each token. As a result, efficiently implementing sparse matrix operations is beyond the capabilities of current PyTorch operators. Typically, previous MoE implementations iterate over each expert and construct the corresponding inputs in a “for loop” style. However, in SMoRA, the large number of experts makes direct for-loop iteration prohibitively time-consuming. Another approach involves using einsum to construct the corresponding LoRA matrices for each token, but this consumes significant GPU memory, making it infeasible for large-scale training and inference. To overcome this, we developed a custom CUDA kernel, indexed_matmul, implemented using TVM, to accelerate sparse matrix computations. Specifically, given two matrices ${\bm{x}}$ and ${\bm{A}}$, we use the index to select the corresponding rows in ${\bm{A}}$ to multiply with ${\bm{x}}$ (while similarly selecting the corresponding columns in ${\bm{B}}$ for computation). This approach enables efficient and flexible sparse matrix multiplication. We evaluated 100 data points to measure the time taken to generate the first token during inference and analyzed the peak GPU memory usage across various methods. As shown in Fig.8, the indexed_matmul method strikes a balance between computational efficiency and memory overhead, significantly outperforming other methods.
Ablation on Activated Rank Number.
In Fig.7, we compare the performance of Vanilla LoRA and SMoRA under different numbers of activated ranks. From the figure, we observe the following: (1) The performance of LoRA improves as the number of total ranks increases; 2) For SMoRA, the best performance is achieved by activating 8 ranks, as the number of activated ranks governs the granularity of information sharing between tasks. In our setup, which consists of 48 tasks divided into 10 task clusters, the differences among tasks can be substantial. Activating too many ranks may force excessive knowledge sharing, potentially leading to task interference. Conversely, using too few shared ranks could result in insufficient utilization of knowledge. Therefore, selecting 8 ranks for activation strikes the optimal balance, effectively managing the trade-off between inter-task information sharing and interference; (3) When SMoRA activates all parameters, it degenerates into Vanilla LoRA, resulting in identical performance. (4) With the same number of activated ranks, SMoRA consistently outperforms LoRA.
@@ 5 Related Work
5.1 PEFT Methods
Fine-tuning large language models (LLMs) requires substantial computational resources, making it impractical in many scenarios. To address this, PEFT methods update only a small subset of parameters while keeping most of the model frozen. Among these, LoRA (Hu et al., 2021) is widely used, reducing trainable parameters by decomposing weight updates into the product of two low-rank matrices, improving efficiency without sacrificing performance. Several extensions of LoRA, such as AdaLoRA (Zhang et al., 2023a), IA3 (Liu et al., 2022), and SoRA (Ding et al., 2023), aim to enhance its flexibility, dynamic rank allocation, and efficiency. However, these methods fix parameters after training, preventing dynamic selection based on input, which limits adaptability in downstream heterogeneous tasks.
@@ 5.1 PEFT Methods
5.2 LoRA with MoE
Some works apply MoE-based methods specifically for multi-task or domain-specific learning. For example, MoELoRA (Liu et al., 2023a) focuses on multi-task learning with explicit task labels in medical applications, while LoRA-MoE (Dou et al., 2023) with soft routing introduces gating mechanisms to allocate LoRA experts during fine-tuning, alleviating catastrophic forgetting. Similarly, SMEAR (Muqeeth et al., 2023) achieves a dynamic fusion of LoRA experts, allowing the model to adjust its parameters adaptively during different tasks. MixLoRA (Li et al., 2024) employs a top-k routing strategy, typically selecting the top-2 LoRA experts for multi-task learning, while HydraLoRA (Tian et al., 2024) proposes an asymmetric structure for its MoE design, further improving parameter utilization in multi-task scenarios. However, these methods rely on coarse-grained expert partitioning, limiting the flexibility of parameter combinations and hindering the model’s ability to dynamically adjust its internal representations. This coarse granularity leads to suboptimal performance, as it fails to fully exploit the potential for more precise and adaptive knowledge sharing across tasks.
In addition to training-time MoE, several post-training methods have also been explored (Zhao et al., 2024a; Wu et al., 2023; Xu et al., 2024). Techniques such as MoLE propose first training individual LoRA modules for each specific task and subsequently training an additional router to control task-specific activation during inference. However, these post-training MoE designs are largely focused on inference-time flexibility and are less relevant to training-time adaptive methods, which are the main focus of this paper.
6 Conclusion
In this paper, we establish a connection between Multi-LoRA MoE and single LoRA by demonstrating that Multi-LoRA can be viewed as a single LoRA with block-wise activation. Through further empirical analysis, we find that finer-grained segmentation and activation of LoRA parameters enhance performance in downstream heterogeneous tasks. Building on these insights, we introduce SMoRA, which incorporates a MoE structure within LoRA to enable fine-grained, rank-wise activation. To broaden SMoRA’s applicability, we employ a training-free load-balancing strategy to effectively distribute load among experts. Additionally, we implement a CUDA kernel using TVM to optimize sparse matrix computations within SMoRA. Experimental results show that SMoRA achieves superior performance in multi-task learning with fewer activated parameters, demonstrating its adaptability and efficiency.
Appendix A Evaluation Benchmark Details
@@ A.1 FLAN-v2 Subset
A.2 Multi-Domain Benchmark
@@ A.4 Evaluation Metrics
Appendix B Main Results on DS-1000
@@
Appendix C Implementation Details of Baseline Methods
In this section, we provide a detailed explanation of the baseline methods used for comparison with the proposed SMoRA. Each method is described with its core idea, mathematical formulation, and specific implementation details. For all methods, we use Kaiming initialization for matrix ${\bm{A}}$ and zero initialization for matrix ${\bm{B}}$ in the LoRA modules, unless otherwise specified.
C.1 LoRA with Different Ranks
@@
Implementation: We test LoRA with ranks of 64, 32, 16, and 8. Matrix ${\bm{A}}$ is initialized using Kaiming initialization, and matrix ${\bm{B}}$ is initialized to zero. This ensures that the initial update $\Delta W=0$, preserving the pre-trained model’s performance at the start of training.
C.2 LoRA MoE with Different Routing Strategies
@@
  • Top-k Gating: For the Top-k routing strategy, we use a routing mechanism similar to Mixtral (Jiang et al., 2024), which selects the top-k experts based on their gating scores and combines their outputs accordingly. Specifically, the gating function is defined as: $g({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}(xW_{g}))$, where only the top-k experts are considered for each input.
@@ C.5 MoSLoRA
Appendix D TVM Implementation of indexed_matmul
TVM (Chen et al., 2018) is an end-to-end deep learning compiler stack designed to optimize machine learning workloads across diverse hardware backends. It enables efficient computation by providing automated tensor optimization techniques such as operation fusion, memory layout transformations, and parallelization, making it particularly suitable for high-performance deep learning tasks.
In our work, we implemented a custom CUDA kernel, indexed_matmul, using TVM to accelerate sparse matrix computations in SMoRA. The key innovation of indexed_matmul lies in its dynamic handling of sparse operations. Instead of performing a full matrix multiplication, the kernel leverages the top-K indices from the gating function to dynamically extract and compute only the relevant rows and columns of the LoRA matrices. This approach significantly reduces computational and memory overhead by avoiding redundant calculations.
@@ Appendix D TVM Implementation of indexed_matmul
  • Efficient Memory Access: The computation was scheduled with fine-grained memory access patterns to minimize memory latency and maximize bandwidth utilization on CUDA-enabled GPUs.
@@
These optimizations enable indexed_matmul to handle large-scale models with higher ranks efficiently, making it both computationally and memory-efficient.
Appendix E Full Results on FLAN-v2
In this section, we present the full results of our experiments conducted on FLAN-v2, covering 10 domains and 48 tasks. The results for the two base models, Llamma-{7b,13b}, are shown in Tab.5 and Tab.6.
@@ Appendix E Full Results on FLAN-v2
Appendix F Difference between SMoRA and MoSLoRA
@@
Appendix G Visualization of Routing Distribution Under Load Balancing Strategy Ablation
@@
Appendix H Comparison of Parameter Counts Across Methods

### v2

@@ 3 Experiments
In Fig.5, we select a specific layer and illustrate the load distribution across all ranks, both with and without the auxiliary load balancing strategy. When the load balancing strategy is applied, the load is evenly distributed across ranks, indicating a relatively balanced workload. In contrast, without the load-balancing strategy, certain ranks experience significantly higher loads, leading to an imbalanced workload distribution. To quantitatively evaluate the load balance, we calculate the MaxVio (maximal violation) value Wang et al. (2024), defined as the difference between the maximum load and the average load, with smaller values indicating a more balanced distribution. With the load balancing strategy, the MaxVio is significantly reduced to 1.23, compared to 5.71 without the strategy. This stark contrast further highlights the effectiveness of the adopted load-balancing strategy in achieving a more uniform load distribution. As shown in the right part of Fig.5, applying the load-balancing strategy results in a 0.34% performance improvement. In Appendix I, we further visualize the routing distribution under the ablation of the load-balancing strategy. The results clearly demonstrate that with the load-balancing strategy removed, the routing becomes unevenly distributed, heavily favoring a subset of experts. This imbalance highlights the critical role of the load-balancing mechanism in ensuring a fair and efficient allocation of tasks across all available experts.
@@
In our approach, the router selects the appropriate ranks to participate in the computation for each token. As a result, efficiently implementing sparse matrix operations is beyond the capabilities of current PyTorch operators. Typically, previous LoRA-MoE implementations iterate over each expert and construct the corresponding inputs in a “for loop" style. However, in SMoRA, the large number of experts makes direct for-loop iteration prohibitively time-consuming. Another approach involves using einsum to construct the intermediate LoRA matrices for each token, but this consumes significant GPU memory, making it infeasible for large-scale training and inference. To overcome this, we developed a custom CUDA kernel, indexed_matmul, implemented using TVM, to accelerate sparse matrix computations. Specifically, given two matrices ${\bm{x}}$ and ${\bm{A}}$, we use the index to select the corresponding rows in ${\bm{A}}$ to multiply with ${\bm{x}}$ (while similarly selecting the corresponding columns in ${\bm{B}}$ for computation). This approach enables efficient and flexible sparse matrix multiplication. We evaluated 100 data points to measure the time taken to generate the first token during inference and analyzed the peak GPU memory usage across various methods. As shown in Fig.6, the indexed_matmul method strikes a balance between computational efficiency and memory overhead, outperforming other implementations.
@@ 4 Related Work
PEFT methods Hu et al. (2021); Zhang et al. (2023a); Liu et al. (2022); Ding et al. (2023) with their low-rank decomposition have substantially reduced the computational cost of adapting LLMs. However, these techniques generally yield static parameters after training, which limits their ability to dynamically adjust to diverse inputs, a crucial feature for heterogeneous downstream tasks.
LoRA-MoE approaches Liu et al. (2023a); Dou et al. (2023); Muqeeth et al. (2023); Li et al. (2024); Tian et al. (2024) typically use routing mechanisms to select or combine different LoRA modules, treating them as experts. However, the coarse-grained expert division can restrict fine-grained, input-specific adjustments, thereby potentially hindering adaptability and optimal performance on diverse tasks. For a more comprehensive discussion, please refer to Appendix B.
5 Conclusion
This paper demonstrated that multi-LoRA MoE can be unified with single LoRA through block-wise activation. Our empirical analysis further revealed that finer-grained rank activation enhances performance on heterogeneous tasks and expands parameter representational capacity. Based on these insights, we introduced SMoRA, which embeds MoE within LoRA for fine-grained, rank-wise activation. Experimental results show that SMoRA achieves superior performance in multi-task learning with fewer activated parameters.
Limitations
One potential limitation of SMoRA’s fine-grained expert partitioning, despite its theoretical FLOPs reduction, pertains to its practical execution efficiency on highly parallel hardware like GPUs. The fine-grained approach leads to highly dispersed matrix computations. While theoretically requiring fewer operations, this fragmentation can result in underutilization of the large compute cores prevalent in modern GPUs, potentially yielding practical speedups that are less pronounced than what theoretical FLOPs might suggest. Conversely, this characteristic could be advantageous for edge devices. The typically smaller compute cores in edge hardware might align better with SMoRA’s fine-grained computational tasks, potentially offering more significant relative efficiency gains in such scenarios.
Appendix A Ablation on Activated Rank Number
In Fig.7, we compare the performance of Vanilla LoRA and SMoRA under different numbers of activated ranks. From the figure, we have the following observations: (1) The performance of LoRA improves as the number of total ranks increases; 2) For SMoRA, the best performance is achieved by activating 8 ranks, as the number of activated ranks governs the granularity of information sharing between tasks. In our setup, which consists of 48 tasks divided into 10 task clusters, the differences among tasks can be substantial. Activating too many ranks may force excessive knowledge sharing, potentially leading to task interference. Conversely, using too few shared ranks could result in insufficient utilization of knowledge. Therefore, selecting 8 ranks for activation strikes the optimal balance, effectively managing the trade-off between inter-task information sharing and interference. (3) When SMoRA activates all parameters, it degenerates into Vanilla LoRA, resulting in identical performance. (4) With the same number of activated ranks, SMoRA consistently outperforms LoRA.
@@ Appendix A Ablation on Activated Rank Number
Appendix B Detailed Related Work
B.1 PEFT Methods
Fine-tuning large language models (LLMs) requires substantial computational resources, making it impractical in many scenarios. To address this, PEFT methods update only a small subset of parameters while keeping most of the model frozen. Among these, LoRA Hu et al. (2021) is widely used, reducing trainable parameters by decomposing weight updates into the product of two low-rank matrices, improving efficiency without sacrificing performance. Several extensions of LoRA, such as AdaLoRA Zhang et al. (2023a), IA3 Liu et al. (2022), and SoRA Ding et al. (2023), aim to enhance its flexibility, dynamic rank allocation, and efficiency. However, these methods fix parameters after training, preventing dynamic selection based on input, which limits adaptability in downstream heterogeneous tasks.
B.2 LoRA with MoE
Some works apply MoE-based methods specifically for multi-task or domain-specific learning. For example, MoELoRA Liu et al. (2023a) focuses on multi-task learning with explicit task labels in medical applications, while LoRA-MoE Dou et al. (2023) with soft routing introduces gating mechanisms to allocate LoRA experts during fine-tuning, alleviating catastrophic forgetting. Similarly, SMEAR Muqeeth et al. (2023) achieves a dynamic fusion of LoRA experts, allowing the model to adjust its parameters adaptively during different tasks. MixLoRA Li et al. (2024) employs a top-k routing strategy, typically selecting the top-2 LoRA experts for multi-task learning, while HydraLoRA Tian et al. (2024) proposes an asymmetric structure for its MoE design, further improving parameter utilization in multi-task scenarios. However, these methods rely on coarse-grained expert partitioning, limiting the flexibility of parameter combinations and hindering the model’s ability to dynamically adjust its internal representations. This coarse granularity leads to suboptimal performance, as it fails to fully exploit the potential for more precise and adaptive knowledge sharing across tasks.
In addition to training-time MoE, several post-training methods have also been explored Zhao et al. (2024a); Wu et al. (2023); Xu et al. (2024). Techniques such as MoLE propose first training individual LoRA modules for each specific task and subsequently training an additional router to control task-specific activation during inference. However, these post-training MoE designs are largely focused on inference-time flexibility and are less relevant to training-time adaptive methods, which are the main focus of this paper.
Appendix C Evaluation Benchmark Details
@@ C.1 FLAN-v2 Subset
C.2 Multi-Domain Benchmark
@@ C.4 Evaluation Metrics
Appendix D Main Results on DS-1000
@@
Appendix E Implementation Details of Baseline Methods
In this section, we provide a detailed explanation of the baseline methods used for comparison with the proposed SMoRA. Each method is described with its core idea, mathematical formulation, and specific implementation details. For all methods, we use Kaiming initialization for matrix ${\bm{A}}$ and zero initialization for matrix ${\bm{B}}$ in the LoRA modules, unless otherwise specified. This ensures that the initial update $\Delta W=B\cdot A=0$, preserving the pre-trained model’s performance at the start of training.
E.1 LoRA with Different Ranks
@@
Implementation: We test LoRA with ranks of 64, 32, 16, and 8. Matrix ${\bm{A}}$ is initialized using Kaiming initialization, and matrix ${\bm{B}}$ is initialized to zero.
E.2 LoRA MoE with Different Routing Strategies
@@
  • Top-k Gating: For the Top-k routing strategy, we use a routing mechanism similar to Mixtral Jiang et al. (2024), which selects the top-k experts based on their gating scores and combines their outputs accordingly. Specifically, the gating function is defined as: $g({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}(xW_{g}))$, where only the top-k experts are considered for each input.
@@ E.5 MoSLoRA
Appendix F TVM Implementation of indexed_matmul
TVM Chen et al. (2018) is an end-to-end deep learning compiler stack designed to optimize machine learning workloads across diverse hardware backends. It enables efficient computation by providing automated tensor optimization techniques such as operation fusion, memory layout transformations, and parallelization, making it particularly suitable for high-performance deep learning tasks.
In our work, we implemented a custom CUDA kernel, indexed_matmul, using TVM to accelerate sparse matrix computations in SMoRA. The key innovation of indexed_matmul lies in its dynamic handling of sparse operations. Instead of performing a full matrix multiplication, the kernel leverages the top-K indices from the gating function to dynamically extract and compute only the relevant rows and columns of the LoRA matrices. This approach reduces computational and memory overhead by avoiding redundant calculations.
@@ Appendix F TVM Implementation of indexed_matmul
Appendix G Full Results on FLAN-v2
In this section, we present the full results of our experiments conducted on FLAN-v2, covering 10 domains and 48 tasks. The results for the two base models, Llama-{7b,13b}, are shown in Tab.5 and Tab.6.
@@ Appendix G Full Results on FLAN-v2
Appendix H Difference between SMoRA and MoSLoRA
@@
Appendix I Visualization of Routing Distribution Under Load Balancing Strategy Ablation
@@
Appendix J Comparison of Parameter Counts Across Methods